# Towards improved property prediction of two-dimensional (2D) materials using many-body Quantum Monte Carlo methods

### v1

@@ Abstract
The field of two-dimensional (2D) materials has grown dramatically in the last two decades. 2D materials can be utilized for a variety of next-generation optoelectronic, spintronic, clean energy, and quantum computation applications. These 2D structures, which are often exfoliated from layered van der Waals (vdW) materials, possess highly inhomogeneous electron densities and can possess short- and long-range electron correlations. The complexities of 2D materials make them challenging to study with standard mean-field electronic structure methods such as density functional theory (DFT), which relies on approximations for the unknown exchange-correlation functional. In order to overcome the limitations of DFT, highly accurate many-body electronic structure approaches such as Diffusion Monte Carlo (DMC) can be utilized. In the past decade, DMC has been used to calculate accurate magnetic, electronic, excitonic, and topological properties in addition to accurately capturing interlayer interactions and cohesion and adsorption energetics of 2D materials. This approach has been applied to 2D systems of wide interest including graphene, phosphorene, MoS2, CrI3, VSe2, GaSe, GeSe, borophene, and several others. In this review article, we highlight some successful recent applications of DMC to 2D systems for improved property predictions beyond standard DFT.
@@ I Introduction
Since the synthesis of graphene in 2004 by Geim and Novoselov that led to the 2010 Nobel Prize [1, 2, 3], there has been an overwhelming interest in the field of two-dimensional (2D) materials. 2D materials are single-layer crystalline structures that have a large lateral dimension compared to their thickness. Oftentimes, these monolayers are exfoliated from layered materials that are held together by weak van der Waals (vdW) bonds. Due to their lack of surface groups or dangling bonds and large surface-to-volume ratio, 2D materials can possess interesting properties that are substantially different from those of their bulk counterparts [1, 2, 3, 4, 5, 6, 7, 8]. In addition, 2D materials can possess enhanced quantum confinement and significantly reduced dielectric screening [7, 8, 9]. These materials also present interesting phenomena such as enhanced carrier mobility[1, 2, 3], reduction of charge carrier scattering [1, 2, 3, 4, 5, 6, 9], superior mechanical properties [10], direct-to-indirect band gap transitions [4], non-trivial topological states [11, 12, 13], and superconductivity [14, 15] and magnetism in 2D [16]. These physical phenomena can be exploited for future applications in optoelectronics, spintronics, quantum computing, and clean energy [17, 18].
@@
In addition to graphene, there exist several other synthesized monoelemental 2D materials such as germanene [19, 20], borophene [21, 22, 23, 24], silicene [25, 26], antimonene [27, 28, 29], phosphorene [30, 31, 32], and tellurene [33, 34]. A popular class of 2D materials for optoelectronic and magnetic applications are transition metal dichalcogenides (TMDs) [4, 5, 6, 35] such as monolayer MoS2 [4, 5, 6], MoSe2 [36, 37], WSe2 [37, 38, 39] and WTe2 [13, 40]. Similarly to TMDs, there exist post-transition metal chalcogenides (PTMCs) [41, 42] such as 2D GaSe [43, 44, 45, 46, 47, 48, 49, 41, 50], GeSe [51, 52] and InSe [53, 54] that have also shown promise for optoelectronic applications. Another class of 2D materials with applications in energy storage are MXenes [55, 56, 57], which are monolayer structures of transition metal carbides, nitrides, or carbonitrides. Monolayer halide-based materials such as CrI3 [16], CrBr3 [58], and NiI2 [59] have also been synthesized as 2D magnets and RuCl3 has been shown to have topological properties [11].
@@
The electronic density in 2D materials is highly inhomogeneous. 2D materials may possess a combination of short-range, strong correlations (if d orbitals are involved) with long-range, weak correlations. These complicated electronic properties of 2D materials make them difficult to model using methods such as Density Functional Theory (DFT) [60, 61]. DFT calculations are, in principle, exact, but rely on approximations for the unknown exchange-correlation (XC) functional that models the many-body quantum interactions [60, 61]. Although modern approximations for density functionals have advanced throughout the last few decades, progressing from local (local density approximation, LDA [62, 63]), to semilocal (generalized gradient approximation, GGA [64, 65] and meta-GGA [66, 67, 68, 69, 70]), to hybrid functionals (i.e., HSE, PBE0 and B3LYP[71, 72, 73, 74, 75, 76, 77, 72]), certain shortcomings remain. The most obvious limitation of DFT is that its results can strongly depend on the XC functional and there is no a priori way to know the significance of the approximations in the chosen functional. In addition, many DFT functionals require explicit corrections to handle vdW interactions [78, 79, 80, 81, 82, 83, 84, 85]. vdW interactions are crucial for low-dimensional layered materials, whose properties are significantly impacted by which vdW correction or functional is used in DFT [86, 87].
@@
Modeling strongly correlated electronic systems poses a challenge within DFT frameworks due to the tendency of standard exchange-correlation functionals (like LDA and GGA) to overly delocalize valence electrons[88]. This issue predominantly stems from the self-interaction error, [62] rendering DFT inadequate for accurately describing strongly localized systems, such as those found in transition metal-based materials containing and electrons. To correct this delocalization error, the Hubbard (U) correction [89], which is an additional onsite Coulomb interaction, can be applied to correlated electronic states. Although self-consistent approaches for obtaining U have been developed [90, 91], the choice of U value can be arbitrary and is often either fit to experiments or selected from previous literature values [92]. The vast majority of interesting 2D materials are electronic, magnetic, and topological materials that contain transition metals with correlated and electrons and DFT+U has been previously applied to handle such correlations [93, 94]. Hybrid/exact-exchange functionals and some meta-GGAs also reduce the degree of self-interaction error, but a consistent high-accuracy treatment remains elusive. The problem of self-interaction error was the main topic of the original LDA DFT paper[62], and it is interesting to note that the problem persists more than 40 years later.
@@
Another major shortcoming of DFT is its underestimation of the band gap for semiconductors and insulators, which results in significant disagreements with experiments. This underestimation is clearly apparent for local and semilocal functionals such as LDA and GGA, with noted improvements for meta-GGA functionals[95, 96, 97]. Significant improvements to the accuracy of DFT electronic structure predictions can be achieved through the use of hybrid functionals, but there are fundamental theoretical issues that can limit DFT’s accuracy[95, 96, 97]. DFT is a ground state theory that maps the electronic structure problem to a fictitious Kohn-Sham system of single electrons occupying energy bands. Therefore, the band gap computed with DFT can be understood as the Kohn-Sham energy gap. For an infinite solid calculated with the exact DFT functional, the DFT gap would be the exact quasiparticle gap [98]. In real materials, electrons feel the effects of surrounding electrons and holes, a phenomenon that is not captured at the quasiparticle level. In order to accurately predict optical excitations (the promotion gap, excitonic effects), these quasi-electron and quasi-hole interactions must be taken into account [99]. These interactions can be obtained using post-DFT methods such as many-body perturbation theory (i.e., the GW approximation) [99, 100, 101] and the Bethe-Salpeter equation (BSE) [99, 102] for electron-hole interactions or by exciting the system within a many-body stochastic theory such as Diffusion Monte Carlo (DMC) [103]. The quasiparticle picture and excitonic effects are extremely important for understanding 2D materials [7, 9]. The reduced dielectric screening in low-dimensional systems can result in strong excitonic effects (exciton binding energies in the 10 - 100 meV range) [7, 8, 9].
@@
In addition to reduced screening, spin-orbit coupling (SOC) can play an important role in understanding the electronic structure of 2D materials [104, 105, 106, 107, 12]. Several TMDs have SOC-induced band splitting in the valence and conduction bands, which can be exploited for spintronic and valleytronic applications. In addition, SOC can induce band inversions and other topological properties in 2D materials [104, 105, 106, 107, 12]. Modeling spin-orbit in 2D materials can be challenging with standard DFT methods. There have been several computational studies that have shown the considerable impact of SOC on the electronic structure and have demonstrated that results can significantly vary depending on the choice of methodology [108, 12, 109, 110, 111, 112, 113].
@@
2D materials are promising and complicated structures that have significant vdW interactions, correlated electronic and magnetic properties, and unique excitonic and topological properties that make them the perfect test bed for more accurate computational methodologies. The strong dependence of the results on the density functional and Hubbard correction can severely limit the reliability of DFT for these 2D systems. In addition, methods that can correct bands gaps and other DFT predictions such as GW and BSE still significantly rely on the Kohn-Sham eigenvalues obtained from a particular DFT functional. For this reason, Diffusion Monte Carlo (DMC) [103] is an ideal method to accurately describe the properties of 2D materials. Although DMC is orders of magnitude more computationally intensive than DFT, it is extremely scalable on modern computers and scales as , where is the number of electrons in the simulation [103]. In addition, the stochastic many-body nature of DMC and its controllable approximations allow for results that have a much weaker dependence on the starting wavefunction and density functional [103].
@@
In this review article, we will give an overview of how the DMC method has recently been applied to a variety of low-dimensional systems. Section II will give a brief overview of the theory, Section III.1 will give a description of how DMC has been applied to correlated 2D magnetic systems, Section III.2 will detail how DMC has been used for improved predictions of electronic properties (including excitonic, spin-orbit, and topological effects), Section III.3 will focus on how DMC has been used to improve the prediction of interlayer interactions, Section III.4 will describe how DMC has been used for the cohesion and adsorption energetics of low-dimensional materials, and Section IV will give concluding remarks and offer future perspectives.
@@ II Theory
In this review article, we survey the application of the DMC [103] method, a real-space quantum Monte Carlo (QMC) method, to low-dimensional systems. Although the exact specifications of the calculations may vary among the works presented in this manuscript, we will provide a broad overview of the theory and approximations used throughout these calculations.
@@
DMC is a projector-based method that can be used to obtain the ground state energy of a many-body system. In this method, the time-dependent Schrödinger equation is recast into the imaginary-time () Schrödinger equation:
@@
where is the Hamiltonian operator which consists of kinetic and potential energy contributions, measures the progress in imaginary time, is the wavefunction at imaginary time , and is the offset of the ground state energy. As , the weight of the DMC configurations with higher energy are damped exponentially. Therefore, the ground state wavefunction is projected out once a steady state has been reached.
@@
The fluctuations in the reweighing process have been shown to be reduced significantly by using a trial (guiding) wavefunction () which transforms the wavefunction in Equation 1 into . The quality of the trial wavefunction can be improved by a set of Jastrow factors [114]:
@@
where represents the Jastrow factors and and represent Slater determinants of up and down spins. The Slater determinant in the trial wavefunction usually comes from a DFT or Hartree Fock (HF) calculation. More general representations constructed from multi-determinant expansions are also possible. In fact, the application of millions of determinants to solids is now possible. [115, 116] The Jastrow factor explicitly includes electron correlations such as parameterized electron-ion, electron-electron, and electron-electron-ion terms. These terms are found through optimization of the wavefunction and the ground state energy and/or variance [117, 118]. The increased accuracy of the trial wavefunctions – being closer to an exact eigenstate – also serves to reduce the number of statistical samples needed to reach a specific error bar, usually reducing computational costs.
@@
DMC formally treats the many-body electron-electron interaction exactly, however it utilizes several approximations to reduce the cost of performing the calculations in the presence of core electrons and the fermion sign problem. Fixed-node DMC produces the ground state energy with the constraint , such that the DMC wavefunction shares the same nodes or phase as the trial wavefunction [63, 119, 120, 103]. This fixed-node approximation is enforced to maintain the antisymmetry of the wavefunction. Given the exact nodal surface of a system from an exact trial wavefunction, the fixed-node DMC method will yield the exact energy. For approximate nodal surfaces, the error introduced is variational (positive) in the energy. Fortunately for approximate trial wavefunctions, this error is typically small, and because the error is variational, different choices of input wavefunction can be tested and the most accurate selected. Practical methods to minimize or test the fixed-node error involve changing the trial wavefunction by adjusting the Hubbard parameter in DFT+U or tuning the amount of exact exchange in hybrid DFT and finding the minimum DMC energy variationally[121, 122, 123, 124, 125]. In addition, there exist more sophisticated techniques to improve the accuracy of the trial wavefunction, such as using multideterminants [126, 116, 115] or backflow transformations [127]. In principle, the accuracy can be systematically improved by using these forms, as well as by optimizing the orbitals entirely within the QMC calculation. This has yet to be done systematically in 2D materials, but several applications in bulk exist[128, 129] and indicate that the approach should be practical.
@@
Another essential aspect of DMC calculations are pseudopotentials. Pseudopotentials are required in DMC to avoid the large cost of performing all-electron calculations. High-quality pseudopotentials are essential for ensuring the accuracy of DMC simulations while simultaneously removing chemically inert core electrons. Luckily, in the past decade, there has been a strong effort to develop DMC-specific norm-conserving pseudopotentials with suitable accuracy [130, 131, 132, 133, 134, 135, 136, 137, 138, 139]. Although the original method for evaluating pseudopotentials in DMC, the locality approximation [140], was non-variational, newer methods such as the T-moves approach [141] restore the variational properties of the overall method.

### v2

# Toward improved property prediction of 2D materials using many-body quantum Monte Carlo methods

@@ Abstract
The field of 2D materials has grown dramatically in the past two decades. 2D materials can be utilized for a variety of next-generation optoelectronic, spintronic, clean energy, and quantum computing applications. These 2D structures, which are often exfoliated from layered van der Waals materials, possess highly inhomogeneous electron densities and can possess short- and long-range electron correlations. The complexities of 2D materials make them challenging to study with standard mean-field electronic structure methods such as density functional theory (DFT), which relies on approximations for the unknown exchange-correlation functional. To overcome the limitations of DFT, highly accurate many-body electronic structure approaches such as diffusion Monte Carlo (DMC) can be utilized. In the past decade, DMC has been used to calculate accurate magnetic, electronic, excitonic, and topological properties in addition to accurately capturing interlayer interactions and cohesion and adsorption energetics of 2D materials. This approach has been applied to 2D systems of wide interest, including graphene, phosphorene, MoS2, CrI3, VSe2, GaSe, GeSe, borophene, and several others. In this review article, we highlight some successful recent applications of DMC to 2D systems for improved property predictions beyond standard DFT.
@@ I Introduction
Since the synthesis of graphene in 2004 by Geim and Novoselov that led to the 2010 Nobel Prize,[1, 2, 3] there has been an overwhelming interest in the field of 2D materials. 2D materials are single-layer crystalline structures that have a large lateral dimension compared to their thickness. Oftentimes, these monolayers are exfoliated from layered materials that are held together by weak van der Waals (vdW) bonds. Due to their lack of surface groups or dangling bonds and large surface-to-volume ratio, 2D materials can possess interesting properties that are substantially different from those of their bulk counterparts.[1, 2, 3, 4, 5, 6, 7, 8] In addition, 2D materials can possess enhanced quantum confinement and significantly reduced dielectric screening.[7, 8, 9] These materials also present interesting phenomena such as enhanced carrier mobility,[1, 2, 3] reduction of charge carrier scattering,[1, 2, 3, 4, 5, 6, 9] superior mechanical properties,[10] direct-to-indirect band gap transitions,[4] nontrivial topological states,[11, 12, 13] and superconductivity [14, 15] and magnetism in 2D.[16] These physical phenomena can be exploited for future applications in optoelectronics, spintronics, quantum computing, and clean energy.[17, 18]
@@
In addition to graphene, there exist several other synthesized monoelemental 2D materials such as germanene, [19, 20] borophene, [21, 22, 23, 24] silicene, [25, 26] antimonene, [27, 28, 29] phosphorene, [30, 31, 32] and tellurene. [33, 34] A popular class of 2D materials for optoelectronic and magnetic applications is transition metal dichalcogenides (TMDs) [4, 5, 6, 35] such as monolayer MoS2, [4, 5, 6] MoSe2, [36, 37] WSe2, [37, 38, 39] and WTe2. [13, 40] Similarly to TMDs, there exist post-transition metal chalcogenides (PTMCs) [41, 42] such as 2D GaSe, [43, 44, 45, 46, 47, 48, 49, 41, 50] GeSe, [51, 52] and InSe [53, 54] that have also shown promise for optoelectronic applications. Another class of 2D materials with applications in energy storage is MXenes, [55, 56, 57] which are monolayer structures of transition metal carbides, nitrides, or carbonitrides. Monolayer halide-based materials such as CrI3, [16] CrBr3, [58] and NiI2 [59] have also been synthesized as 2D magnets, and RuCl3 has been shown to have topological properties. [11]
@@
The electronic density in 2D materials is highly inhomogeneous. 2D materials may possess a combination of short-range, strong correlations (if d orbitals are involved) with long-range, weak correlations. These complicated electronic properties of 2D materials make them difficult to model using methods such as density functional theory (DFT).[60, 61] DFT calculations are, in principle, exact but rely on approximations for the unknown exchange-correlation (XC) functional that models the many-body quantum interactions.[60, 61] Although modern approximations for density functionals have advanced throughout the past few decades, progressing from local (local density approximation [LDA] [62, 63]), to semilocal (generalized gradient approximation [GGA] [64, 65] and meta-GGA [66, 67, 68, 69, 70]), to hybrid functionals (i.e., HSE, PBE0, and B3LYP[71, 72, 73, 74, 75, 76, 77, 72]), certain shortcomings remain. The most obvious limitation of DFT is that its results can strongly depend on the XC functional, and there is no a priori way to know the significance of the approximations in the chosen functional. In addition, many DFT functionals require explicit corrections to handle vdW interactions.[78, 79, 80, 81, 82, 83, 84, 85] The vdW interactions are crucial for low-dimensional layered materials, whose properties are significantly impacted by which vdW correction or functional is used in DFT.[86, 87]
@@
Modeling strongly correlated electronic systems poses a challenge within DFT frameworks due to the tendency of standard XC functionals (like LDA and GGA) to overly delocalize valence electrons.[88] This issue predominantly stems from the self-interaction error,[62] rendering DFT inadequate for accurately describing strongly localized systems, such as those found in transition metal–based materials containing and electrons. To correct this delocalization error, the Hubbard (U) correction, [89] which is an additional on-site Coulomb interaction, can be applied to correlated electronic states. Although self-consistent approaches for obtaining U have been developed, [90, 91] the choice of U value can be arbitrary and is often either fit to experiments or selected from previous literature values.[92] The vast majority of interesting 2D materials are electronic, magnetic, and topological materials that contain transition metals with correlated and electrons, and DFT+U has been previously applied to handle such correlations.[93, 94] Hybrid/exact-exchange functionals and some meta-GGAs also reduce the degree of self-interaction error, but a consistent high-accuracy treatment remains elusive. The problem of self-interaction error was the main topic of the original LDA DFT paper,[62] and the problem persists more than 40 years later.
@@
Another major shortcoming of DFT is its underestimation of the band gap for semiconductors and insulators, which results in significant disagreements with experiments. This underestimation is clearly apparent for local and semilocal functionals such as LDA and GGA, with noted improvements for meta-GGA functionals.[95, 96, 97] Significant improvements to the accuracy of DFT electronic structure predictions can be achieved through the use of hybrid functionals, but there are fundamental theoretical issues that can limit DFT’s accuracy.[95, 96, 97] DFT is a ground-state theory that maps the electronic structure problem to a fictitious Kohn-Sham system of single electrons occupying energy bands. Therefore, the band gap computed with DFT can be understood as the Kohn-Sham energy gap. For an infinite solid calculated with the exact DFT functional, the DFT gap would be the exact quasiparticle gap.[98] In real materials, electrons feel the effects of surrounding electrons and holes, a phenomenon that is not captured at the quasiparticle level. To accurately predict optical excitations (e.g., the promotion gap, excitonic effects), these quasi-electron and quasi-hole interactions must be taken into account.[99] These interactions can be obtained by using post-DFT methods such as many-body perturbation theory (i.e., the GW approximation) [99, 100, 101] and the Bethe-Salpeter equation (BSE) [99, 102] for electron-hole interactions or by exciting the system within a many-body stochastic theory such as diffusion Monte Carlo (DMC).[103] The quasiparticle picture and excitonic effects are extremely important for understanding 2D materials.[7, 9] The reduced dielectric screening in low-dimensional systems can result in strong excitonic effects (exciton binding energies in the 10–100 meV range).[7, 8, 9]
@@
In addition to reduced screening, spin-orbit coupling (SOC) can play an important role in understanding the electronic structure of 2D materials.[104, 105, 106, 107, 12] Several TMDs have SOC-induced band splitting in the valence and conduction bands, which can be exploited for spintronic and valleytronic applications. In addition, SOC can induce band inversions and other topological properties in 2D materials.[104, 105, 106, 107, 12] Modeling spin-orbit in 2D materials can be challenging with standard DFT methods. There have been several computational studies that have shown the considerable impact of SOC on the electronic structure and have demonstrated that results can significantly vary depending on the choice of methodology.[108, 12, 109, 110, 111, 112, 113]
@@
2D materials are promising and complicated structures that have significant vdW interactions, correlated electronic and magnetic properties, and unique excitonic and topological properties that make them the perfect test bed for more accurate computational methodologies. The strong dependence of the results on the density functional and Hubbard correction can severely limit the reliability of DFT for these 2D systems. In addition, methods that can correct bands gaps and other DFT predictions such as GW and BSE still significantly rely on the Kohn-Sham eigenvalues obtained from a particular DFT functional. For this reason, DMC [103] is an ideal method to accurately describe the properties of 2D materials. Although DMC is orders of magnitude more computationally intensive than DFT, it is extremely scalable on modern computers and scales as , where is the number of electrons in the simulation.[103] In addition, the stochastic many-body nature of DMC and its controllable approximations allow for results that have a much weaker dependence on the starting wavefunction and density functional.[103]
Although there are major shortcomings of DFT that can be overcome by using many-body methods such as DMC, DFT is an absolutely essential tool in the electronic structure community. For a majority of problems in physics, chemistry, and materials science, standard DFT is suitable to obtain important qualitative trends in predicted properties. In fact, for closed-shell systems and scenarios without strong correlations, DFT does an adequate job of describing electronic properties, and systematic trends are well described. The low computational expense also allows more systems to be studied than with many-body methods, facilitating the construction of materials databases. However, fundamentally DFT and many-body approaches such as QMC have different properties, so we consider it best to ask when, where and how each method may be most productively applied. Unlike DFT, the majority of QMC methods offer variational properties enabling the few approximations in them to be systematically tested and potentially reduced in small enough systems. QMC methods (such as DMC) can describe strong electron correlations and van der Waals interactions more accurately than DFT, making them very suitable for 2D materials. They are therefore ideal for as-yet unsynthesized novel materials, where there is no empirical data to guide DFT approximation selection or where there are known issues with DFT approximations for the particular materials class. For the materials covered in this review, QMC can be applied directly, but can also be used to validate the selection of exchange-correlation functional.
The systematic trends from DFT functionals can be used to our advantage when attempting to understand complex phenomena. For example, the underbinding and subsequent lattice constant errors of GGA were used as a method to identify new exfoliable vdW materials. [87] One might ask to what extent DFT is successful and when it should be trusted as a complement to DMC. In turn, one might ask in what situations is DMC absolutely necessary. Due to the fact that DFT is significantly less computationally expensive than DMC allows us to perform DFT calculations at a much larger scale, in terms of number of calculations and system sizes. Fundamentally, however, the methods have different properties, so it is better to ask when and where each method may be best applied. Most importantly, we can use our DMC results as a metric to compare various DFT calculations to, and find which DFT functional best matches the DMC results. This can aid in the design of new DFT functionals (i.e., improving existing vdW functionals). It also allows us to make an informed decision about which DFT functional (or value of Hubbard U, amount of HF mixing in hybrid functionals, which vdW correction, etc.) to use for a system of interest in order to obtain more complicated properties that are difficult to obtain with DMC (i.e., band structure, magnetic anisotropy, simulations at larger system sizes). This is especially relevant for systems that do not have an existing experimental benchmark to compare DFT to or if we are focusing on properties that are not easily obtained by experiments (i.e., interlayer binding energy, magnetic exchange). DFT and DMC generally focus on ground state properties, since they are both based on a procedure of energy minimization. Despite this fact, time-dependent DFT [114] can be applicable to a broad range of excited states while many-body approaches such as DMC can yield accurate excited state results (i.e., excitonic effects). While modern QMC rests on the foundations of the development work done in the 1980s and 1990s, critical advances made since then include the development of twist averaging (enabling practical calculation of metallic materials), more accurate trial wavefunction optimization methods, including the critical breakthrough of the linear method/energy minimization, and the development of better methods for pseudopotentials, which are critical for the heavier elements in dichalcogenides. These are cited at several points in the review [115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136].
@@
In this review article, we provide an overview of how the DMC method has recently been applied to a variety of low-dimensional systems. Section II provides a brief overview of the theory, Section III.1 describes how DMC has been applied to correlated 2D magnetic systems, Section III.2 details how DMC has been used for improved predictions of electronic properties (including excitonic, spin-orbit, and topological effects), Section III.3 focuses on how DMC has been used to improve the prediction of interlayer interactions, Section III.4 describes how DMC has been used for the cohesion and adsorption energetics of low-dimensional materials, and Section IV provides concluding remarks and offers future perspectives.
@@ II Theory
In this review article, we survey the application of the DMC [103] method, a real-space quantum Monte Carlo (QMC) method, to low-dimensional systems. Although the exact specifications of the calculations may vary among the works presented in this manuscript, we provide a broad overview of the theory and approximations used throughout these calculations.
@@
DMC is a projector-based method that can be used to obtain the ground-state energy of a many-body system. In this method, the time-dependent Schrödinger equation is recast into the imaginary-time () Schrödinger equation:
@@
where is the Hamiltonian operator, which consists of kinetic and potential energy contributions; measures the progress in imaginary time; is the wavefunction at imaginary time ; and is the offset of the ground-state energy. As , the weight of the DMC configurations with higher energy is damped exponentially. Therefore, the ground-state wavefunction is projected out once a steady state has been reached.
@@
The fluctuations in the reweighting process (adjusting the statistical weight of sampled configurations to account for differences between the sampled and targeted distributions) have been shown to be reduced significantly by using a trial (guiding) wavefunction (), which transforms the wavefunction in Equation 1 into . The quality of the trial wavefunction can be improved by a set of Jastrow factors: [137]
@@
where represents the Jastrow factors; , represent the -th Slater determinants of up and down spins in the multideterminant expansion; and represents the weight of the -th determinant configuration. The Slater determinants in the trial wavefunction usually come from a DFT or Hartree-Fock (HF) calculation. For most of the work presented in this review, a single-determinant trial wavefunction is used. More general representations constructed from multideterminant expansions are also possible. In fact, the application of millions of determinants to solids is now possible.[138, 139] The Jastrow factor explicitly includes electron correlations such as parameterized electron-ion, electron-electron, and electron-electron-ion terms. These terms are found through optimization of the wavefunction and the ground-state energy and/or variance.[115, 116] The increased accuracy of the trial wavefunctions—being closer to an exact eigenstate—also serves to reduce the number of statistical samples needed to reach a specific error bar, usually reducing computational costs.
@@
DMC formally treats the many-body electron-electron interaction exactly; however, it utilizes several approximations to reduce the cost of performing the calculations in the presence of core electrons and the fermion sign problem. Fixed-node DMC produces the ground-state energy with the constraint , such that the DMC wavefunction shares the same nodes or phase as the trial wavefunction.[63, 140, 141, 103] This fixed-node approximation is enforced to maintain the antisymmetry of the wavefunction. Given the exact nodal surface of a system from an exact trial wavefunction, the fixed-node DMC method will yield the exact energy. For approximate nodal surfaces, the error introduced is variational (positive) in the energy. Fortunately for approximate trial wavefunctions, this error is typically small, and because the error is variational, different choices of input wavefunction can be tested and the most accurate selected. There have been some guided efforts to estimate the fixed-node error [142, 143, 144, 145, 117], and particularly for low-dimensional materials [146, 147, 118, 148], but the number of results is small enough that it is premature to infer general trends. One approach to estimate the error involves performing higher accuracy quantum chemistry calculations such as selected Configuration Interaction (sCI). These can be used either directly or as a source of improved trial wavefunction in QMC, enabling a sensitivity and convergence analysis of the fixed-node error provided a sufficient number of determinants can be considered. It has been found that the magnitude of this error is entirely system dependent. It is important to note that sCI cannot always serve as a direct benchmark, unless the relevant space of determinants is not very large. However, sCI provides a way to estimate DMC nodal errors in systems where the sCI energy can be sufficiently converged in terms of basis set and system size (e.g., in primitive cells of solids and in small molecules). For large enough systems, sCI becomes impractically expensive to converge, and the DMC errors cannot be directly estimated via sCI. For example, the fixed-node error of Si systems was found to be extremely negligible for the ground state (1.3(2)), [144] using the recovery of the nearly-exact correlation energy as an error metric. This fixed-node error has been estimated to be similar in magnitude for systems containing Sn (2), [147] Ru (4), [118] and Cl (3) [118] but significantly higher for systems containing Tb (12) [147] and Mn (9).[147] However, it has been demonstrated that these large biases can cancel when calculating quantities such as molecular binding energies.[147] It is important to note that systems for which one cannot compute well converged sCI energies (due to system size), it may still be possible to compute well converged CC energies. CC (ideally, CC calculations with single, double, and perturbative triple excitations (CCSD(T)) in the complete basis set limit) may be used to estimate the fixed-node error in systems where the CC ansatz works well, such as in closed-shell systems and systems that are largely single-reference. For example, the fixed-node bias was estimated for phosphorene from a fully converged CCSD(T) calculation of a P4 cluster [149]. It has been demonstrated that CC and DMC have strong agreement for ground state energy difference-based quantities for large noncovalent complexes [143] and materials dominated by vdW interactions.[142] There have been some assessments of the fixed node errors for single determinant DMC band gaps, where it has been reported that for certain systems, the fixed node error for the ground state and the excited state are different, leading to a slight bias in the gap. [144, 149, 148] On the other hand, there have been instances where the fixed node error for the single determinant DMC band gap has been negligible.[146] Practical methods to minimize or test the fixed-node bias involve changing the trial wavefunction by adjusting the Hubbard parameter in DFT+U or varying the amount of exact exchange in hybrid DFT and finding the minimum DMC energy variationally.[150, 118, 151, 152, 153] In cases where the fixed-node error is high, there exist more sophisticated techniques to improve the accuracy of the trial wavefunction, such as optimizing the orbitals within QMC [119, 154, 155], using multideterminants [156, 139, 138] or backflow transformations.[157, 145] Backflow transformations can be particularly challenging and computationally expensive for solids.[145, 148] It has been reported that the gain in accuracy from backflow is overshadowed by errors from finite-size (FS) and nonlocal pseudopotenitals.[145, 148, 157, 158] Extensive benchmarking of backflow for molecules and solids can be found in Ref. 145, and benchmarking of backflow for a 2D system (hexagonal boron nitride [hBN]) can be found in Ref. 148. Accuracy can also be improved by optimizing the orbitals entirely within the QMC calculation. This has yet to be done systematically in solid-state materials for production-level calculations (there have yet to be any examples for 2D materials), but preliminary applications indicate that the approach should be practical [119, 154, 155] and it has been used in studies of cuprates [153, 159, 160], the Ce volume collapse [161], and for phonon studies. [162, 163]
In addition to variational Monte Carlo (VMC) and DMC, other robust techniques such as Coupled Cluster (CC) and Auxiliary Field QMC (AFQMC) have been refined and applied to difficult problems throughout the last decade. Although these methods have not yet been extensively applied to 2D or quasi-2D systems (as is the case for VMC and DMC), CC has been used to compute the electronic spectrum of TMDs such as monolayer MoS2, MoSe2, WS2, and WSe2,[164] and AFQMC has been applied to study Moiré systems, including a metal-insulator transition in a semiconductor heterobilayer model.[165] Aside from low-dimensional materials, CC has been used to obtain the quasiparticle band structure of simple bulk solids [166, 167] and AFQMC has been used for a wider range of materials including oxides, diamond, LiF and Al. [168, 169, 170] We expect the expanded application to 2D materials to follow in the coming years.
@@
Pseudopotentials are an essential aspect of DMC calculations. Pseudopotentials are required in DMC to avoid the large cost of performing all-electron calculations. High-quality pseudopotentials are essential for ensuring the accuracy of DMC simulations while simultaneously removing chemically inert core electrons. Luckily, in the past decade, there has been a strong effort to develop DMC-specific norm-conserving pseudopotentials with suitable accuracy.[120, 121, 122, 123, 124, 125, 126, 127, 128, 129] Although the original method for evaluating pseudopotentials in DMC, the locality approximation, [130] was nonvariational, newer methods such as the T-moves approach [131] restore the variational properties of the overall method.
The accuracy of DMC results is also impacted by FS errors.[103] Specifically, there are two types of FS errors that can impact the DMC simulations. The first type, one-body FS errors, arises from insufficient sampling of the Brillouin zone.[103] This error can be significantly reduced by using twist-averaged boundary conditions. Various twist-averaging (TA) schemes have been applied to solids throughout the years, ranging from the use of the computationally inexpensive Baldereschi-point grids [171, 132, 133, 134] to more recent enhanced TA techniques for magnetic metals.[135] In most common production-level calculations of 2D materials, the total energy is converged at the DFT level with respect to -point grid size, and this grid is used as the subsequent number of twists in the DMC calculations. Ideally, one would wish to converge the number of twists at the DMC level, but this is a computationally expensive endeavor. The second type, two-body FS errors, arises from the artificial interaction of periodic XC holes.[103] This can be minimized by performing the simulation at multiple supercell sizes and then extrapolating to the thermodynamic (infinite-size) limit. This extrapolation is essential for achieving accurate observable quantities, and careful calculations at supercells large enough to eliminate this bias must be conducted. In addition to FS errors, errors can arise from the choice of time step . In DMC, the walkers are propagated in imaginary time by statistically sampling a Green’s function. This relies on the Green’s function projector within the short time approximation, which becomes exact as approaches 0. For this reason, finite time steps can introduce errors in the DMC energies, and careful analysis and testing must be performed when choosing a time step to minimize the target uncertainty.[103] In most of 2D material works presented in this review, generously small time steps on the order of 0.01 Ha-1 (0.27 eV-1) were chosen to minimize time-step error.[149, 172, 173, 109, 174, 110, 175, 146]
For systems with significant spin-orbit effects, the QMC methods have been generalized to handle electron spins as quantum variables so that any spin state, including noncollinear spins, can be described. The first key difference from spinless Hamiltonians is that one-particle spatial-only orbitals are replaced by two-component spinors. The spinors can be written as , where orbitals are calculated in spinor-based self-consistent approaches (DFT, Dirac-Fock, etc.). The spinors also include spin variables that enter the spin functions , while are complex constants such that . The fixed-node Slater-Jastrow trial functions are generalized to accommodate the spinors as given by
For example, we can see that a singlet wavefunction in Eq. (2) is built from determinants of matrices, while in the spinor wavefunction, there are determinants of matrices (because spins vary, the position space does not factorize into up- and down-spin subspaces). Correspondingly, the fixed-node approximation is generalized to the fixed-phase approximation because the trial function is inherently complex. We can write and substitute it into the imaginary-time Schrödinger equation. Decomposition into real and imaginary parts leads to an exact equation for the amplitude given by
The fixed-phase approximation imposes the unknown phase to be equal to the phase of the trial state as given by
The fixed-phase approximation is variational because the repulsive potential can only raise the energy for any approximate . The spin-orbit terms are typically included in the pseudopotentials with the formalism generalized accordingly; see further details in Refs. 136, 176.